6.The company called ALOR is a commercial firm. It sells three products A, B and C. The forecasts data for the next accounting period are the following: UNIT SALES SELLING | ACQUISITION |COMMERCIAL PRODUCTS (units) A1 PRICE (E/unit) 0,000 6,7504.050 5,625 5,625 COST (E/unit) COST (%/revenues) 10% 4% 2% 8,000 20,000 2,587.5 3,262.5 We know that the fixed costs amount to a total of 78.000,000 of which 22,0000,000 are traceable and direct to product A and 12,000,000 direct traceable relate to product B, and 5,000,000 are traceable and direct to product C, while the remaining costs are common fixed costs for the company Required: 1. Prepare the Segmented Income Statement Report (Showing each product and the whole company) 2. Calculate the global breakeven point and break it down for each product in Revenues. 3. Assuming that the sales mix doesnt change, calculate the total revenues (Showing each product and the whole company) needed to obtain a target operating income of e 19,500,000 4. Calculate the critical points of each product and, based on this, decide if it would be convenient to continue commercialising the three products. Justify your answer. 5. Assuming that the total revenues provided at the required number 1 doesnt change (€225,000,000), the % of variable cost doesnt change, the total fixed cost remain equal and the new sales mix is 30%, 30% and 40% for A, B and C, respectively. Prepare the new Segmented Income Statement Report (Showing each product and the whole company) and comment the impact this fact has on operating income. Note: If you find it convenient, you can express calculations in millions of euros.
Answer #1
1.. A B C Total
Unit sales 10000 8000 20000
Unit selling price 6750 5625 5625
Sales $ 67500000 45000000 112500000 225000000
Less: Acquistion costs 40500000 20700000 65250000 126450000
Less: Commercial cost 6750000 1800000 2250000 10800000
Contribution 20250000 22500000 45000000 87750000
Less:Traceable fixed costs 22000000 12000000 5000000 39000000
Segmented Income -1750000 10500000 40000000 48750000
Less: Common fixed costs 39000000
Net Income 9750000
Workings for Global break-even point:
2.. A B C
1.Unit sales 10000 8000 20000
Per unit
2.Selling price 6750 5625 5625
3.Acquisition cost 4050 2587.5 3262.5
4.Commercial cost(%*S.P) 675 225 112.5
5.Marginal contribution /unit(2-(3+4)) 2025 2812.5 2250
6.Total contribution(5*1) 20250000 22500000 45000000 87750000
7.Traceable Fixed costs 22000000 12000000 5000000 39000000
8.Segmented net income(6-7) -1750000 10500000 40000000 48750000
9.Common fixed costs 39000000
10.Net Income(8-9) 9750000
A B C Total
1.Unit sales 10000 8000 20000 38000
Per unit
2.Selling price 6750 5625 5625
3.Acquisition cost 4050 2587.5 3262.5
4.Commercial cost(%*S.P) 675 225 112.5
5.Marginal contribution /unit(2-(3+4)) 2025 2812.5 2250
6. Sales mix %( Sales units/Total sales units) 26% 21% 53%
7. 5*6 533 592 1184
8. Wt. av. Contribution (Sum of Row. 7) 2309
Global BEP=Total fixed costs/Wt. av. Contribution per unit
78000000/2309=
33781
units
No.of units at the Global Break-even point:
Sales mix %( Sales units/Total sales units) 26% 21% 53%
1.Sales Units= Global BEP*Sales mix % 8783 7094 17904
2.Selling price/unit 6750 5625 5625
Total sales/revenues $ at BEP(1*2) 59285405 39903638 100709181 199898224
3..
Target Sales in units=(Fixed costs+Target operating income)/Wt. av. Cont. Margin *Sales mix %
Target Sales in units (A)=(78000000+19500000)/2309 *26 %= 10979 units
Target Sales in units (B)=(78000000+19500000)/2309 *21 %= 8867 units
Target Sales in units (C)=(78000000+19500000)/2309 *53%= 22380 units
Target revenues
A-- 10979*6750= 74108250
B-- 8867*5625= 49876875
C--22380*5625= 125887500
Total revenues 249872625
Target opg. Income-19500000 A B C Total
Unit sales 10979 8867 22380
Unit selling price 6750 5625 5625
Sales $ 74108250 49876875 125887500 249872625
Acquistion costs 44464950 22943363 73014750 140423063
Commercial cost 7410825 1995075 2517750 11923650
Contribution 22232475 24938438 50355000 97525913
Traceable fixed costs 22000000 12000000 5000000 39000000
Segmented Income 232475 12938438 45355000 58525913
Common fixed costs 39000000
Net Income(subject to Rounding-off error) 19525913
4. Critical point of the 3 products:
A B C
Marginal contribution/unit(as in 2.) 2025 2812.5 2250
Traceable Fixed costs 22000000 12000000 5000000
Individual BEP/Critical point(in units)--Fixed costs/Marginal contribution 10864 4267 2222
Product B & C 's critical/BE point are well below the current sales volume & hence are contributing to their own as well as overall fixed costs.
Whereas,
Product A 's Critical point is above the current sales level --hence it is not meeting its own fixed costs , leave alone contributing to overall fixed costs.
So, the company will benefit by commercialising more of B & C.
5.... A B C Total
Unit sales 11400 11400 15200 38000
Unit selling price 6750 5625 5625
Sales $ 76950000 64125000 85500000 226575000
Acquistion costs 46170000 29497500 49590000 125257500
Commercial cost 7695000 2565000 1710000 11970000
Contribution 23085000 32062500 34200000 89347500
Traceable fixed costs 22000000 12000000 5000000 39000000
Segmented Income 1085000 20062500 29200000 50347500
Common fixed costs 39000000
Net Income 11347500
Impact on Operating Income:
Net Income increases by 1597500
Segmented income increases by 2835000 9562500 -10800000
Per-unit workings for 5.. Above A B C
1.Unit sales 11400 11400 15200 38000
Per unit
2.Selling price 6750 5625 5625
3.Acquisition cost 4050 2587.5 3262.5
4.Commercial cost(%*S.P) 675 225 112.5
5.Marginal contribution /unit(2-(3+4)) 2025 2812.5 2250
6.Total contribution(5*1) 23085000 32062500 34200000 89347500
7.Traceable Fixed costs 22000000 12000000 5000000 39000000
8.Segmented net income(6-7) 1085000 20062500 29200000 50347500
9.Common fixed costs 39000000
10.Net Income(8-9) 11347500